# thanks: Supported by the Natural Science Foundation of Shanghai (No. 23ZR1429300) and the Innovation Funds of CNNC (Lingchuang Fund, contract No. CNNC-LCKY-202234)

### v1

@@ Abstract
Reactor physics is the study of neutron properties, focusing on using models to examine the interactions between neutrons and materials in nuclear reactors. Artificial intelligence (AI) has made significant contributions to reactor physics, e.g., in operational simulations, safety design, real-time monitoring, core management and maintenance. This paper presents a comprehensive review of AI approaches in reactor physics, especially considering the category of Machine Learning (ML), with the aim of describing the application scenarios, frontier topics, unsolved challenges and future research directions. From equation solving and state parameter prediction to nuclear industry applications, this paper provides a step-by-step overview of ML methods applied to steady-state, transient and combustion problems. Most literature works achieve industry-demanded models by enhancing the efficiency of deterministic methods or correcting uncertainty methods, which leads to successful applications. However, research on ML methods in reactor physics is somewhat fragmented, and the ability to generalize models needs to be strengthened. Progress is still possible, especially in addressing theoretical challenges and enhancing industrial applications such as building surrogate models and digital twins.
@@ I Introduction
Nowadays, most of the literature may not directly refer to the study of "reactor physics", but refers to the terms "neutronics", "nuclear reactor theory" or "reactor analysis", which have almost similar meanings bib:5 , with the commonality that the relevant methods or research ideas are used, or that the physics related to the neutrons of the fission reaction is involved, which is also the method of definition in this review. Frequent research objectives in the literature are the study of the determination of the neutron position in reactor cores, the reaction of neutrons with the surrounding medium, and the study of the spatio-temporal distribution of neutrons, including the improvement of efficiency by ML methods. In addition, the synthesis and organization of the literature related to neutronic interactions with matter in nuclear reactors by AI methods is still fragmented. For numerical simulation techniques in the field of reactor physics, the latest publication bib:6 classified the studies into six categories: (i) nuclear data processing and resonance calculation models, (ii) cross-section homogenization techniques, (iii) steady-state and transient neutron transport methods, (iv) Monte Carlo approaches and applications, (v) nuclear reactor design and analysis, and (vi) methods for sensitivity and uncertainty analysis. Inspired by this research, we will investigate the field of reactor physics, collate the works of literature focusing on frontier applications of methods in ML to the field, and demonstrate what reactor physics is, how it can be solved with the help of ML methods to solve problems related to it, and how to enhance and improve the practical efficiency of the methods.
@@ I.2 What the Artificial Intelligence Methods are in Literature
In recent years, ML methods, particularly neural networks (NNs) and their variants, have made significant contributions to the advancement of nuclear reactor physics. However, it is crucial to recognize that the application of AI in this field is not limited to NNs alone. Other static learning methods, such as support vector machines (SVM), random forests (RF), and support vector regression (SVR), also play pivotal roles in addressing various challenges in reactor physics.
Whether in the simulation of nuclear reactor operation, the design of the reactor core, or the real-time monitoring, management, and maintenance of nuclear power plants, AI methods have been widely applied. The current advanced methods are already present in the industrial applications of nuclear power plants, but still suffer from the problem of mutual constraints of accuracy and time scale. As powerful autonomous learners, the emergence of AI methods, including but not limited to NNs, precisely presents a better solution to these problems.
Figure 2: The chart shows the histogram of literature statistics based on the Scopus platform. The trend indicates a significant increase in the number of publications from 2020 onwards. The data is collected from the earliest year available up to 2023, with the horizontal axis representing the years and the vertical axis indicating the number of publications.
@@
Figure 3: Refined statistics of ML methods in the literature include several labels categorized by algorithmic applications, including CNN, DNN, RNN, Reinforcement Learning (RL), Transfer Learning (TL), Long and Short-Term Memory (LSTM), and Fully Convolutional Networks (FCN), etc., in Deep Learning (DL), and there is also a further class of neural networks based on deep learning frameworks, generally referred to as PINN bib:49 , which is an efficient and feasible approach for solving equations and is particularly useful for solving the equations governing the reactions.
@@ I.3 What is this Review About
Based on the analysis above, in this survey, we focus on how AI can be used to address different problems of reactor physics, e.g., solving the neutron transport equation, the building blocks of reactor physics, aspects related to transport theory, the toolset that is already available, future directions and recent trends, as well as issues related to sensitivity and uncertainty analysis. Depending on these problems, we show how high-fidelity nuclear reactor simulations can be customized in literature through parameter prediction, data augmentation, and the use of transfer learning.
This review aims to cover most of the frontier ML-related articles related to nuclear reactor physics, and to learn about advances in machine learning in nuclear physics, see this article bib:69a . A practical scenario of the ML approach to reactor physics in the article is shown in Fig. 5. We stress that the objective of this work is not to compare the performance of ML methods since they were developed for different goals and to address different problems. Thus, this paper can serve as a comprehensive guide for navigating these fast-growing ML-related techniques and methodologies in reactor physics. The primary research question is to determine what reactor physics is and the ML-based applications to the simulation, design, and safety regulation of nuclear reactors. In addition, the review considers researches that address a broader topic origin from but more than reactor physics, which includes important theories of nuclear reactors such as thermal hydraulics.
@@ II Fundamentals of Reactor Physics Theory
In multiscale modeling, energy levels are often discretized to represent different physical scales or interaction regimes within the reactor core. If neutrons in the reactor core are assumed to have isotropic properties within a specific reaction zone, the directional component of neutron flight can be neglected. This leads to the simplification of the model into two categories: steady-state and transient equations, which are further elaborated in Section 3 of this paper. The steady-state approach focuses on reactor behavior over extended periods, examining the reactor’s lifetime and its final state, without taking individual neutron reactions into account. This removes the correlation between the neutron flux and time, as governed by the neutron diffusion equation. On the other hand, transient equations, such as the point kinetic equations (PKEs), focus on the time evolution of the system. In this case, the spatial position of individual neutrons is irrelevant; instead, the neutron population and the reaction rates are modeled as a function of time, describing the overall kinetics of the reactor rather than the movement of individual neutrons.
@@ II.1 Governing Equations
In equation (1), $\varphi(\vec{r},\vec{\omega},E,t)$ denotes the neutron flux, which depends on the position vector $\vec{r}$, energy $E$, and direction of motion $\vec{\omega}$ at time $t$. This flux is a fundamental parameter in nuclear reactor physics, as it describes the distribution and intensity of neutrons in the reactor. The term $\varphi^{\prime}$ refers to the neutron flux at energy $E^{\prime}$ and direction $\vec{\omega}^{\prime}$, representing contributions from neutrons scattering into the specified state $(\vec{r},\vec{\omega},E,t)$ from other energies and directions. The neutron transport equation, in its most general form, involves seven independent variables: three spatial coordinates in $\vec{r}$, two angular coordinates in $\vec{\omega}$, energy $E$, and time $t$ bib:9 . This comprehensive model, though accurate, is often simplified in practical reactor design through approximations such as the diffusion equation.
Equation (1) can describe the neutron transport behavior in a medium precisely. More specifically, in the transport process, neutrons will interact with the medium and undergo reactions such as scattering and absorption, and these will affect the neutron flux. In the equation, $\Sigma$ denotes the macroscopic cross-section, i.e., the probability that a certain process (e.g., collision, fission, absorption) occurs for a neutron, and other notions are standard.
@@
While point kinetics equations (PKEs) have been successfully employed in reactors of all sizes throughout nuclear history, it is known that the underlying diffusion theory can exhibit significant errors in regions with large gradients of neutron current (flux), particularly at reactor boundaries bib:14 . There are several numerical methods for solving this equations, including finite element methods (FEM), finite volume methods (FVM), and finite difference methods (FDM) bib:34 . However, the accuracy of these methods is essentially limited by the number of nodes and geometry of the problem bib:61 .
On the other hand, for the case where the spatial dependence of the neutron flux is negligible, reduced-order models consisting of PKEs have been developed bib:39 . PKEs consist of a system of coupled nonlinear stiff differential equations that model the kinetics of reactor variables. This is a reduced-order model of the Boltzmann neutron transport equation and the Bateman equation describing the 3D spatial and temporal kinetics of a nuclear reactor bib:47 . The PKEs are derived under the approximation of neglecting the shape of the neutron flux and the neutron density distribution bib:47 . As the basis of diffusion theory, Fick’s law is essentially a P1 approximation to the neutron transport equation, which is also known as the telegraph equation, further neglecting the derivative term of the neutron flux density concerning time, and therefore unable to accurately characterize the actual neutron transport process in some sense bib:62 .
PKEs, first deduced by Henry bib:63 , reduce the mathematical problem from a system of integral differential equations to a system of ordinary differential equations bib:64 , which allows for the rapid solution of time-dependent neutron transport problems. The equation is generally written in 1D and has the following basic form bib:65 .
@@
In this paper, we plan to integrate the extant mainstream AI techniques for solutions to important governing equations in reactor physics. Table 1 contains a synthesis of the kind of AI approaches that have been synthesized in the field of reactor physics, as well as the literature in which the approach has been used.
Recently, with the development of AI technology, the use of AI, especially deep learning (DL), to solve complex differential equations is becoming a hot research topic in the field of numerical computation bib:66 . The advantage of resorting to DL instead of traditional neural networks is that it can solve the problem of high computational cost for problems with high-dimensional physical variables and can handle complex boundary conditions bib:40 . Based on artificial neural networks (ANNs), in literature, DL methods have been proposed to solve the neutron diffusion problem for continuous neutron flux distributions without the need for prior regional discretization bib:34 .
In the process of a neutron reaction, the fissile isotopes in the fuel are depleted continuously due to neutron-induced reactions and radioactive decay, and will also change with the reaction bib:67 . To calculate the changes in the composition within the fuel during reactor operation accurately, it is necessary to establish a function of the change in each of the isotopes involved in the reaction over time, that is, the burnup equation. The burnup equation is generally written in the form
@@ III.2.5 Thermo-hydraulic coupling
The nuclear reactor is a complex system with a wide range of physical fields such as neutron density field, flow field, temperature field, etc. widely present in the reactor system bib:110a . There are interactions between these physical fields, so there is a strong coupling relationship between them. And these coupling relations are not negligible during reactor operation, especially during transient processes. Additionally, For neutronics, the neutron fluence rate determines the distribution of fission power inside the core, which affects the temperature field of the fuel rods and ultimately changes the whole temperature field. As for thermal hydraulics, the thermodynamic parameters strongly affect the macroscopic cross section. The interaction of these physical fields forms the basis for the coupling of neutronics and thermo-hydraulics bib:110b . Therefore, we will broadly introduce here the machine learning applications in the scenario of reactor thermal hydrodynamics coupled with reactor physics.
@@
In addition, Aghili Nasr et al. bib:110y tackled a core challenge in coupled simulations: dynamically updating neutron cross sections as a function of evolving state variables. Their method employs Gaussian Process Regression (GPR) to substitute traditional logarithmic correlations for temperature-dependent cross section updates. Implemented within a multiphysics framework based on OpenFOAM, the GPR approach was validated on benchmark problems and a 2D axisymmetric Molten Salt Reactor configuration. The results demonstrate that, particularly for isotopes with prominent absorption resonances (e.g., $\mathrm{{}^{232}Th}$ and $\mathrm{{}^{238}U}$), the method significantly influences the effective multiplication factor ($k_{\text{eff}}$) and transient peak power predictions—highlighting the importance of robust coupling for reactor safety analysis.
@@ IV.2.3 Applications in Safety Design
Core optimization aims to adjust key state parameters of a reactor core—such as power distribution, temperature field, and reactivity coefficients—to ensure both safe operation and optimal performance. Typically, two major classes of optimization methods are employed: intelligent optimization algorithms and gradient descent–based methods.
@@ Acknowledgments
This research was partially sponsored by the Natural Science Foundation of Shanghai (no. 23ZR1429300, 22ZR1421900), the Innovation Funds of CNNC (Lingchuang Fund, contract no. CNNC-LCKY-202234), and the National Natural Science Foundation of China (grant no.12471377).

### v2

@@ Abstract
Reactor physics is the study of neutron properties, focusing on using models to examine the interactions between neutrons and materials in nuclear reactors. Artificial intelligence (AI) has made significant contributions to reactor physics, e.g., in operational simulations, safety design, real-time monitoring, core management and maintenance. This paper presents a comprehensive review of AI approaches in reactor physics, especially considering the category of Machine Learning (ML, that we named also AI/ML to recall AI name we found in articles), with the aim of describing the application scenarios, frontier topics, unsolved challenges and future research directions. From equation solving and state parameter prediction to nuclear industry applications, this paper provides a step-by-step overview of ML methods applied to steady-state, transient and combustion problems. Most literature works achieve industry-demanded models by enhancing the efficiency of deterministic methods or correcting uncertainty methods, which leads to successful applications. However, research on ML methods in reactor physics is somewhat fragmented, and the ability to generalize models needs to be strengthened. Progress is still possible, especially in addressing theoretical challenges and enhancing industrial applications such as building surrogate models and digital twins.
@@ I Introduction
Nowadays, most of the literature may not directly refer to the study of "reactor physics", but refers to the terms "neutronics", "nuclear reactor theory" or "reactor analysis", which have almost similar meanings bib:5 , with the commonality that the relevant methods or research ideas are used, or that the physics related to the neutrons of the fission reaction is involved, which is also the method of definition in this review. Frequent research objectives in the literature are the study of the determination of the neutron position in reactor cores, the reaction of neutrons with the surrounding medium, and the study of the spatio-temporal distribution of neutrons, including the improvement of efficiency by ML methods. In addition, the synthesis and organization of the literature related to neutronic interactions with matter in nuclear reactors by AI/ML methods is still fragmented. For numerical simulation techniques in the field of reactor physics, the latest publication bib:6 classified the studies into six categories: (i) nuclear data processing and resonance calculation models, (ii) cross-section homogenization techniques, (iii) steady-state and transient neutron transport methods, (iv) Monte Carlo approaches and applications, (v) nuclear reactor design and analysis, and (vi) methods for sensitivity and uncertainty analysis. Inspired by this research, we will investigate the field of reactor physics, collate the works of literature focusing on frontier applications of methods in ML to the field, and demonstrate what reactor physics is, how it can be solved with the help of ML methods to solve problems related to it, and how to enhance and improve the practical efficiency of the methods.
@@ I.2 What the Artificial Intelligence Methods are in Literature
In recent years, ML methods, particularly neural networks (NNs) and their variants, have made significant contributions to the advancement of nuclear reactor physics. However, it is crucial to recognize that the application of AI/ML in this field is not limited to NNs alone. Other static learning methods, such as support vector machines (SVM), random forests (RF), and support vector regression (SVR), also play pivotal roles in addressing various challenges in reactor physics.
Whether in the simulation of nuclear reactor operation, the design of the reactor core, or the real-time monitoring, management, and maintenance of nuclear power plants, AI/ML methods have been widely applied. The current advanced methods are already present in the industrial applications of nuclear power plants, but still suffer from the problem of mutual constraints of accuracy and time scale. As powerful autonomous learners, the emergence of AI/ML methods, including but not limited to NNs, precisely presents a better solution to these problems.
Figure 2: Literature Publication Trends in Reactor Physics: This histogram illustrates the annual publication count from the earliest available data through 2023, highlighting a notable surge in research output since 2020. The horizontal axis denotes the years, and the vertical axis shows the number of publications, offering insights into the growing academic interest in this field.
@@
Figure 3: Application of AI/ML Methods in Reactor Physics: A Comparative Analysis. This bar chart categorizes and compares the frequency of various AI/ML methodologies in the literature, including CNN, DNN, RNN, Reinforcement Learnin (RL), Transfer Learning (TL), Long and Short-Term Memor (LSTM), and Fully Convolutional Networks (FCN) within Deep Learning (DL), as well as the specialized PINN bib:49 approach. The x-axis lists the ML methods, and the y-axis indicates the corresponding number of publications, reflecting the diverse algorithmic applications in reactor physics research.
@@ I.3 What is this Review About
Based on the analysis above, in this survey, we focus on how AI/ML can be used to address different problems of reactor physics, e.g., solving the neutron transport equation, the building blocks of reactor physics, aspects related to transport theory, the toolset that is already available, future directions and recent trends, as well as issues related to sensitivity and uncertainty analysis. Depending on these problems, we show how high-fidelity nuclear reactor simulations can be customized in literature through parameter prediction, data augmentation, and the use of transfer learning.
This review aims to cover most of the frontier ML-related articles related to nuclear reactor physics, and to learn about advances in machine learning in nuclear physics, refer to the review bib:69a . A practical scenario of the ML approach to reactor physics in the article is shown in Fig. 5. We stress that the objective of this work is not to compare the performance of ML methods since they were developed for different goals and to address different problems. Thus, this paper can serve as a comprehensive guide for navigating these fast-growing ML-related techniques and methodologies in reactor physics. The primary research question is to determine what reactor physics is and the ML-based applications to the simulation, design, and safety regulation of nuclear reactors. In addition, the review considers researches that address a broader topic origin from but more than reactor physics, which includes important theories of nuclear reactors such as thermal hydraulics.
@@ II Fundamentals of Reactor Physics Theory
In multiscale problems, energy is typically divided into different hierarchical levels. If neutrons in the reactor core are assumed to have isotropic properties within a specific reaction zone, the directional component of neutron flight can be neglected. This leads to the simplification of the model into two categories: steady-state and transient equations, which are further elaborated in Section 3 of this paper. The steady-state approach focuses on reactor behavior over extended periods, examining the reactor’s lifetime and its final state, without taking individual neutron reactions into account. This removes the correlation between the neutron flux and time, as governed by the neutron diffusion equation. On the other hand, transient equations, such as the point kinetic equations (PKEs), focus on the time evolution of the system. In this case, the spatial position of individual neutrons is irrelevant; instead, the neutron population and the reaction rates are modeled as a function of time, describing the overall kinetics of the reactor rather than the movement of individual neutrons.
@@ II.1 Governing Equations
In equation (1), $\Phi(\vec{r},\vec{\omega},E,t)$ denotes the neutron flux, which depends on the position vector $\vec{r}$, energy $E$, and direction of motion $\vec{\omega}$ at time $t$. This flux is a fundamental parameter in nuclear reactor physics, as it describes the distribution and intensity of neutrons in the reactor. The term $\Phi^{\prime}$ refers to the neutron angular flux at energy $E^{\prime}$ and direction $\vec{\omega}^{\prime}$, representing contributions from neutrons scattering into the specified state $(\vec{r},\vec{\omega},E,t)$ from other energies and directions. The neutron transport equation, in its most general form, involves seven independent variables: three spatial coordinates in $\vec{r}$, two angular coordinates in $\vec{\omega}$, energy $E$, and time $t$ bib:9 . This comprehensive model, though accurate, is often simplified in practical reactor design through approximations such as the diffusion equation.
Equation (1) can describe the neutron transport behavior in a medium precisely. More specifically, in the transport process, neutrons will interact with the medium and undergo reactions such as scattering and absorption, and these will affect the neutron flux. In the equation, $\Sigma$ denotes the macroscopic cross-section, i.e., the probability that a certain process occurs for a neutron, for example, collision, which includes scattering (either elastic and inelastic scattering), and absorption (either capture, or fission, or others). Other notions are standard.
@@
While neutron diffusion equation have been successfully employed in reactors of various sizes, their application can lead to significant errors due to the underlying diffusion theory. These errors arise from several factors, including the diffusion approximation itself, spatially integral representations, and simplifications in energy cross-section interactions. In particular, regions with large gradients of neutron current (flux), especially near reactor boundaries, can amplify these errors. Such spatial approximations and their associated inaccuracies are well-established features of the method, rather than unforeseen limitationsbib:14 ; P1 . There are several numerical methods for solving this equations, including finite element methods (FEM), finite volume methods (FVM), and finite difference methods (FDM) bib:34 . However, the accuracy of these methods is essentially limited by the number of nodes and geometry of the problem bib:61 .
As the basis of diffusion theory, Fick’s law is essentially related to the $P_{1}$ approximation of the time-dependent neutron transport equation (the approximation is also known as the neutron telegraph equation), and the diffusion equation further neglects the derivative term of the neutron flux density with respect to time. Since the parabolic nature of the diffusion equation predicts that the particles will have infinite velocities, the particles in the tails of the distribution function will appear at infinite distance from the source in time $t=0+$, but the neutron telegraph equation has finite particle velocities, the value of which is incorrectly P1 . Thus in some sense it cannot accurately characterize the actual neutron transport processbib:62 .
On the other hand, for the case where the spatial dependence of the neutron flux is negligible, reduced-order models consisting of PKEs have been developed bib:39 . PKEs consist of a system of coupled nonlinear stiff differential equations that model the kinetics of reactor variables. This is a reduced-order model of the Boltzmann neutron transport equation and the Bateman equation describing the 3D spatial and temporal kinetics of a nuclear reactor bib:47 . The PKEs are derived under the approximation of neglecting the shape of the neutron flux and the neutron density distribution bib:47 .
PKEs, first deduced by Henry bib:63 , reduce the mathematical problem from a system of integral differential equations to a system of ordinary differential equations bib:64 , which allows for the rapid solution of time-dependent neutron transport problems. The equation is generally written in 0D and has the following basic form bib:65 .
@@
In this paper, we plan to integrate the extented mainstream ML techniques for solutions to important governing equations in reactor physics. Table 1 contains a synthesis of the kind of AI/ML approaches that have been synthesized in the field of reactor physics, as well as the literature in which the approach has been used.
Recently, with the development of ML technology, the use of ML, especially deep learning (DL), to solve complex differential equations is becoming a hot research topic in the field of numerical computation bib:66 . The advantage of resorting to DL instead of traditional neural networks is that it can solve the problem of high computational cost for problems with high-dimensional physical variables and can handle complex boundary conditions bib:40 . Based on artificial neural networks (ANNs), in literature, DL methods have been proposed to solve the neutron diffusion problem for continuous neutron flux distributions without the need for prior regional discretization bib:34 .
In the process of a neutron reaction, the fissile isotopes in the fuel are depleted continuously due to neutron-induced reactions and radioactive decay, and will also change with the reaction bib:67 . To calculate the changes in the composition within the fuel during reactor operation accurately, it is necessary to establish a function of the change in each of the isotopes involved in the reaction over time, that is, the burnup equation. The burnup equation is generally written in the form of equation (4).
@@ III.2.5 Thermo-hydraulic coupling
The nuclear reactor is a complex system with a wide range of physical fields such as neutron density field, flow field, temperature field, etc. widely present in the reactor system bib:110a . There are interactions between these physical fields, so there is a strong coupling relationship between them. And these coupling relations are not negligible during reactor operation, especially during transient processes. Additionally, for neutronics, the neutron fluence rate determines the distribution of fission power inside the core, which affects the temperature field of the fuel rods in addition to the one of water, and ultimately changes the whole temperature field. As for thermal hydraulics, the thermodynamic parameters strongly affect the macroscopic cross section. The interaction of these physical fields forms the basis for the coupling of neutronics and thermo-hydraulics bib:110b . Therefore, we will broadly introduce here the machine learning applications in the scenario of reactor thermal hydrodynamics coupled with reactor physics.
@@
In addition, Aghili Nasr et al. bib:110y tackled a core challenge in coupled simulations: dynamically updating neutron cross sections as a function of evolving state variables. Their method employs Gaussian Process Regression (GPR) to substitute traditional logarithmic correlations for temperature-dependent cross section updates. Implemented within a multiphysics framework based on OpenFOAM, the GPR approach was validated on benchmark problems and a 2D axisymmetric Molten Salt Reactor configuration. The results demonstrate that, particularly for isotopes with prominent absorption resonances (e.g., $\mathrm{{}^{232}Th}$ and $\mathrm{{}^{238}U}$), the method significantly influences the effective multiplication factor ($k_{eff}$) and transient peak power predictions—highlighting the importance of robust coupling for reactor safety analysis.
@@ IV.2.3 Applications in Safety Design
Core optimization aims to adjust key state parameters of a reactor core—such as power distribution, temperature field, and reactivity coefficients—to ensure both safe operation and optimal performance. Typically, two major classes of optimization methods are employed: gradient descent–based methods and evolutionary algorithms (EA).
@@ Acknowledgments
This research was partially sponsored by the Natural Science Foundation of Shanghai (no. 23ZR1429300), the Innovation Funds of CNNC (Lingchuang Fund, contract no. CNNC-LCKY-202234), and the National Natural Science Foundation of China (grant no.12471377).